# OpenScan: A Benchmark for Generalized Open-Vocabulary 3D Scene Understanding

### v2

@@ 4.2 The Impact of Pre-trained Vocabulary Size
To study the impact of the pre-trained vocabulary size on the GOV-3D task, we conduct experiments based on RegionPLC [37] for 3D semantic segmentation. Figure 4 reports the mIoU and mAcc scores under different pre-training vocabulary sizes, . Results show that the majority of the linguistic aspects of object attributes do not exhibit a notable enhancement as increases, reflected by both mIoU and mAcc scores, aligning with our expectations. Some linguistic aspects of object attributes show relatively low performances and exhibit random jitters. Among the eight linguistic aspects, the aspect material illustrates an enhancement in mIoU and a marginal improvement in mAcc as increases. This improvement can be attributed to the framework adopted by RegionPLC, which associates 3D objects with language through explicit visual image captioning models, providing detailed descriptions of visual attributes like material and color for each 3D object. Therefore, as the vocabulary size increases, more objects are processed by the image captioning model to produce visual descriptions that ultimately improve the semantic segmentation results for the aspect material.
@@
This observation suggests that simply increasing the size of the object vocabulary during training may not effectively enhance the generalization capability of OV-3D models. This limitation can be attributed to existing OV-3D benchmarks, like ScanNet [8], ScanNet200 [30], and ScanNet++ [40], which primarily focus on object classes and lack object-related attributes. While increasing the size of the object vocabulary during training can help improve the performance in the OV-3D task, as demonstrated by the results from PLA [10] and RegionPLC [37], this approach is not suitable for the more challenging GOV-3D task. The significant performance gap between the two tasks cannot be resolved simply by transferring the OV-3D technique into the GOV-3D task.
4.3 The Impact of the Query Form
@@
In the benchmark annotation process, we adopt a query generation step to associate attributes and object classes. An ideal query should contain an attribute name and the relation knowledge between the attribute and the corresponding object class. We report the effect of employing a query template (e.g., “this term is made of wood”) versus not using a query template (e.g., “wood”) in the GOV-3D task, as shown in Table 6. We evaluate the 3D instance segmentation results of OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24] models using different query templates. We note that, as expected, the performances of these models improve when using the query template, reflected in metrics in terms of AP, AP50, and AP25. SAI3D and MaskClustering appear to be sensitive to different query templates, while OpenMask3D and Open3DIS exhibit greater robustness across different query templates.
@@
The observed results can be attributed to the fact that current large-scale vision-language models (VLMs), such as CLIP [27], face challenges in classifying object attributes when minor commonsense knowledge is needed in our GOV-3D task, as stated in [39]. Given that most of the existing OV-3D models rely on VLMs like CLIP [27] for open-vocabulary comprehension, they inherit this limitation of VLMs for lacking in commonsense. Therefore, by incorporating query templates containing the relation between attribute names and the corresponding object classes as commonsense knowledge, the performances of OV-3D models have improved in the GOV-3D task.
4.4 Qualitative Results
@@
We present qualitative results from Open3DIS [24] on our OpenScan benchmark. We evaluate Open3DIS across eight linguistic aspects, as shown in Figure 5. It demonstrates that Open3DIS can comprehend specific linguistic aspects such as synonym and material. When exploring the affordance aspect by querying “keep food cold” for the target object, Open3DIS can successfully identify the “refrigerator” as the target object but struggles to generate a correct 3D mask. Additionally, Open3DIS fails to generate predictions for other linguistic aspects. These observations align with the quantitative results in Table 4.
4.5 Failure Cases Analysis
@@
Figure 6 shows a failure case of Open3DIS [24] when applied to GOV-3D, despite its strong performance on OV-3D. While Open3DIS correctly identifies the object class (e.g., “piano”), it fails to recognize the associated object attribute (e.g., “this term has 88 keys”). To investigate this discrepancy, we analyze the CLIP [27] image-text similarity scores in Open3DIS, given that most OV-3D models rely on VLMs like CLIP [27] for 3D predictions. Our analysis reveals that CLIP assigns lower image-text similarity scores to the object attribute compared to the object class, suggesting that its intrinsic attribute knowledge is limited. This observation demonstrates that GOV-3D presents greater challenges compared to OV-3D. A promising direction for GOV-3D involves integrating attribute knowledge into VLMs like CLIP, which may improve image-text alignment and enable more reliable predictions in GOV-3D.
5 Conclusion
@@
In this paper, we address the constraints of the classic Open-Vocabulary 3D Scene Understanding (OV-3D) task, which is limited in handling object attributes beyond object classes. We introduce a more challenging task, called Generalized Open-Vocabulary 3D Scene Understanding (GOV-3D), to comprehensively evaluate the generalization capability of OV-3D models. To facilitate research on the GOV-3D task, we construct a large-scale benchmark named OpenScan, which consists of 341 attributes across 8 linguistic aspects. We systematically evaluate the latest OV-3D models on the OpenScan benchmark, revealing their challenges in understanding attributes beyond object classes. We also conduct experiments to investigate the impact of the pre-training vocabulary size and query form, demonstrating that the generalization ability can be enhanced by utilizing query templates rather than scaling up the vocabulary size during training. We further explore a promising direction for the GOV-3D task by integrating attribute knowledge into VLMs of the OV-3D models. We believe that our OpenScan benchmark can facilitate future research on improving the generalization capability of OV-3D models.
@@
Appendix A Web Interface for Manual Annotation
We implement a web interface for manual annotation of the visual linguistic aspect (e.g., material), as shown in Figure 7. Annotators are shown an interactive 3D mesh of a scene, a list of target objects, and a list of attributes. Users can control the 3D mesh from different viewpoints interactively by rotating, zooming in, zooming out, and panning to observe the scene from various viewpoints. When users select a 3D mesh by clicking the mouse in the scene, the target object will be highlighted and the corresponding object ID and object class will be displayed. The annotation process requires annotators to first select a target object in the 3D mesh (e.g., table of ID 2) and then select a primary attribute that belongs to the target object (e.g., stone). Finally, annotators click the confirm button to submit and store the annotations. Once the selected objects are annotated and confirmed, the corresponding object in the object list will show a check mark symbol. Additionally, to address visual ambiguity issues in 3D object appearance, our annotation process allows users to review the scene’s video sequences, providing contextual visual cues to resolve uncertainties about target objects’ visual attributes during annotation.
@@ Appendix A Web Interface for Manual Annotation
| OpenMask3D [34] | 7.7 | 8.2 | 8.7 | 14.2 | 16.3 | 12.0 | 10.0 | 14.7 | 9.7 |
| SAI3D [41] | 4.0 | 5.5 | 7.4 | 10.6 | 8.6 | 8.8 | 7.3 | 8.3 | 6.7 |
| MaskClustering [36] | 6.0 | 6.2 | 4.1 | 8.0 | 11.7 | 10.3 | 9.3 | 8.1 | 6.8 |
| Open3DIS [24] | 11.5 | 17.6 | 14.5 | 18.5 | 27.7 | 18.6 | 16.6 | 23.4 | 15.6 |
@@
| OpenMask3D [34] | 9.6 | 10.7 | 11.2 | 18.1 | 18.4 | 15.2 | 12.8 | 17.8 | 12.2 |
| SAI3D [41] | 6.4 | 8.2 | 10.6 | 14.9 | 13.0 | 13.8 | 11.5 | 13.3 | 10.1 |
| MaskClustering [36] | 10.3 | 11.5 | 7.2 | 13.5 | 23.1 | 18.7 | 16.1 | 14.0 | 12.0 |
| Open3DIS [24] | 14.3 | 22.2 | 18.3 | 22.6 | 32.9 | 23.1 | 19.7 | 28.5 | 19.2 |
@@
| OpenMask3D [34] | 10.8 | 12.4 | 13.9 | 20.4 | 18.9 | 18.0 | 14.4 | 20.5 | 14.1 |
| SAI3D [41] | 8.5 | 11.1 | 13.1 | 18.6 | 15.9 | 17.5 | 14.7 | 18.6 | 12.8 |
| MaskClustering [36] | 12.3 | 13.5 | 9.4 | 16.9 | 25.1 | 24.1 | 19.0 | 18.1 | 14.6 |
| Open3DIS [24] | 16.2 | 24.0 | 20.2 | 24.8 | 35.9 | 24.9 | 22.7 | 31.9 | 21.3 |
Appendix B Implementation Details
In this section, we report implementation details of the OV-3D models [34, 41, 36, 24, 25, 10, 37] in our experiments as follows:
OpenMask3D. In the class-agnostic mask proposal module, we employ the Mask3D [31] architecture trained on the ScanNet200 [30] training set. For 2D mask proposals, we use SAM [18] with ViT-H as the backbone. We utilize the pre-trained CLIP [27] visual encoder of ViT-L/14 at a 336 pixel resolution to extract image features with 768 dimensions. We set the number of queries to 150, following the implementation of OpenMask3D [34] and Mask3D [31], to ensure a sufficient number of mask proposals for the GOV-3D task.
SAI3D. We employ Semantic-SAM [19] with Swin-L as the backbone to generate 2D mask proposals. The number of queries is set to 150 to ensure sufficient mask proposals for the GOV-3D task.
MaskClustering. We utilize CropFormer [26] as a 2D mask predictor. For 2D mask proposals, we use CLIP [27] visual encoder of ViT-H/14 to extract image features. We follow MaskClustering [36] to adopt the post-processing approach from OVIR-3D [22] to refine the output 3D instances. Specifically, we employ the DBSCAN [11] algorithm to partition disconnected point clusters.
Open3DIS. We utilize the class-agnostic 3D proposal network ISBNet [23] trained on the ScanNet200 [30] training set as 3D proposal. We employ the 2D-Guided-3D Instance Proposal Module in Open3DIS [24]. For 2D mask proposals, we adopt Grounded-SAM [29] as 2D segmentor, which incorporates a Swin-T-based Grounding-DINO [21] decoder and SAM [18] with ViT-H as the backbone.
OpenScene. We employ OpenSeg [13] for image feature extraction and a 2D-3D ensemble model in OpenScene [25]. We utilize MinkowskiNet18A [6] as the 3D backbone during 3D distillation.
PLA. We utilize a model trained on the ScanNet [8] partition of B15/N4, where B15/N4 indicates 15 base and 4 novel categories. We adopt a SparseUNet16 architecture based on sparse convolutions UNet [14] as our 3D encoder for semantic segmentation and integrate the CLIP [27] text encoder as the final classifier.
RegionPLC. We utilize a model trained on the ScanNet [8] partition of B15/N4, where B15/N4 represents 15 base and 4 novel categories. We employ a sparse-convolution-based UNet [14] of SparseUNet16 as the 3D encoder for semantic segmentation, leveraging the CLIP [27] text encoder as the final classifier.
@@ Appendix B Implementation Details
| OpenMask3D [34] | 19.1 | 12.7 | 28.0 | 26.5 | 9.1 | 0.1 | 41.8 | 16.9 | 23.1 | 10.7 | 18.8 |
| SAI3D [41] | 13.1 | 10.3 | 19.2 | 5.6 | 6.5 | 0.1 | 19.7 | 12.8 | 14.9 | 10.4 | 11.3 |
| MaskClustering [36] | 12.8 | 18.7 | 26.3 | 10.4 | 6.3 | 0.3 | 25.1 | 13.2 | 4.8 | 3.5 | 12.1 |
| Open3DIS [24] | 32.9 | 27.1 | 35.4 | 33.0 | 24.6 | 2.4 | 43.1 | 33.8 | 29.9 | 20.9 | 28.3 |
@@
| OpenMask3D [34] | 23.9 | 16.0 | 30.4 | 30.3 | 11.6 | 0.1 | 44.5 | 20.0 | 29.7 | 15.2 | 22.1 |
| SAI3D [41] | 19.6 | 15.7 | 26.9 | 9.5 | 10.2 | 0.1 | 31.8 | 18.4 | 22.6 | 16.2 | 17.1 |
| MaskClustering [36] | 23.8 | 31.6 | 41.0 | 19.7 | 12.0 | 0.5 | 37.0 | 23.2 | 9.3 | 7.6 | 20.6 |
| Open3DIS [24] | 40.9 | 32.7 | 39.0 | 38.6 | 30.7 | 3.5 | 46.3 | 38.8 | 37.6 | 27.8 | 33.6 |
@@
| OpenMask3D [34] | 27.4 | 19.1 | 32.5 | 33.5 | 13.3 | 0.1 | 47.0 | 21.7 | 34.8 | 21.4 | 25.0 |
| SAI3D [41] | 26.4 | 20.5 | 32.2 | 18.5 | 13.7 | 0.2 | 41.0 | 23.4 | 31.4 | 22.0 | 22.9 |
| MaskClustering [36] | 31.3 | 38.4 | 49.7 | 23.0 | 16.7 | 0.6 | 41.6 | 28.2 | 16.1 | 10.9 | 25.6 |
| Open3DIS [24] | 44.7 | 35.3 | 42.6 | 42.2 | 33.5 | 5.1 | 48.3 | 42.4 | 41.8 | 31.6 | 36.7 |

### v3

@@ Introduction
To study the impact of the pre-trained vocabulary size (i.e., the number of pre-trained object classes) on the GOV-3D task, we conduct experiments based on RegionPLC (Yang et al. 2024) for 3D semantic segmentation. Figure 4 reports the mIoU and mAcc scores under different pre-training vocabulary sizes , on the ScanNet (Dai et al. 2017) and ScanNet200 (Rozenberszki et al. 2022) datasets. Results show that the majority of the linguistic aspects of object attributes do not exhibit a notable enhancement as increases, reflected by both mIoU and mAcc scores, aligning with our expectations. Some linguistic aspects of object attributes show relatively low performances and exhibit random jitters. Among the eight linguistic aspects, the aspect material illustrates an enhancement in mIoU and a marginal improvement in mAcc as increases. This improvement can be attributed to the framework adopted by RegionPLC, which associates 3D objects with language through explicit visual image captioning models, providing detailed descriptions of visual attributes like material and color for each 3D object. Therefore, as the vocabulary size increases, more objects are processed by the image captioning model to produce visual descriptions that ultimately improve the semantic segmentation results for the aspect material.
@@
This observation suggests that simply increasing the size of the object vocabulary during training may not effectively enhance the generalization capability of OV-3D models. This limitation can be attributed to existing OV-3D benchmarks, like ScanNet (Dai et al. 2017), ScanNet200 (Rozenberszki et al. 2022), and ScanNet++ (Yeshwanth et al. 2023), which primarily focus on object classes and lack object-related attributes. While increasing the size of the object vocabulary during training can improves the OV-3D performance, as demonstrated by the results from PLA (Ding et al. 2023) and RegionPLC (Yang et al. 2024), this approach is not suitable for the more challenging GOV-3D task. The significant performance gap between the two tasks cannot be resolved simply by transferring the OV-3D technique into GOV-3D.
The Impact of the Query Form
@@
In benchmark annotation, we generate queries linking attributes to object classes. An ideal query should contain an attribute name and the relation between the attribute and the corresponding object class. Table 5 shows the effect of using a query template (e.g., “this term is made of wood”) versus a plain term (e.g., “wood”) in GOV-3D. We evaluate the 3D instance segmentation of OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024), reporting the mean score of all attributes in OpenScan. Note that, as expected, using the query template improves model performance, as shown by higher AP, AP50, and AP25. SAI3D and MaskClustering are more sensitive to query templates, while OpenMask3D and Open3DIS show greater robustness.
@@
These results stem from the fact that VLMs, like CLIP (Radford et al. 2021), struggle with attribute classification in GOV-3D when minor commonsense knowledge is required, as stated in (Ye et al. 2023). Since most OV-3D models rely on VLMs like CLIP (Radford et al. 2021) for open-vocabulary comprehension, they inherit VLMs’ commonsense limitations. Thus, incorporating query templates that link attributes to object classes as commonsense knowledge improves OV-3D models’ performances in GOV-3D.
Qualitative Results
@@
We present qualitative results from Open3DIS (Nguyen et al. 2024) on our OpenScan benchmark. We evaluate Open3DIS across eight linguistic aspects, as shown in Figure 5. It demonstrates that Open3DIS can comprehend specific linguistic aspects such as synonym and material. When exploring the affordance aspect by querying “keep food cold” for the target object, Open3DIS can successfully identify the “refrigerator” as the target object but struggles to generate a correct 3D mask. Additionally, Open3DIS fails to generate predictions for other linguistic aspects. These observations align with the quantitative results in Table 3.
Failure Cases Analysis
@@
Figure 6 shows a failure case of Open3DIS (Nguyen et al. 2024) when applied to GOV-3D, despite its strong performance on OV-3D. While Open3DIS correctly identifies the object class (e.g., “piano”), it fails to recognize the associated object attribute (e.g., “this term has 88 keys”). To investigate this discrepancy, we analyze the CLIP (Radford et al. 2021) image-text similarity scores in Open3DIS, given that most OV-3D models rely on VLMs like CLIP (Radford et al. 2021) for 3D predictions. Our analysis reveals that CLIP assigns lower image-text similarity scores to the object attribute compared to the object class, suggesting that its intrinsic attribute knowledge is limited. This observation demonstrates that GOV-3D presents greater challenges compared to OV-3D. A promising direction for GOV-3D involves integrating attribute knowledge into VLMs like CLIP, which may improve image-text alignment and enable more reliable predictions.
Conclusion
@@
In this paper, we address the constraints of the classic Open-Vocabulary 3D Scene Understanding (OV-3D) task, which is limited in handling object attributes beyond object classes. We introduce a more challenging task, called Generalized Open-Vocabulary 3D Scene Understanding (GOV-3D), to comprehensively evaluate the generalization capability of OV-3D models. To facilitate research on the GOV-3D task, we construct a large-scale benchmark named OpenScan, which consists of 341 attributes across 8 linguistic aspects. We systematically evaluate the OV-3D models on the OpenScan benchmark, revealing their challenges in understanding attributes beyond object classes. We also conduct experiments to investigate the impact of the pre-trained vocabulary size and query form, demonstrating that the generalization ability can be enhanced by utilizing query templates rather than scaling up the vocabulary size during training. We further explore a promising direction for the GOV-3D task by integrating attribute knowledge into VLMs of the OV-3D models. We believe our OpenScan benchmark can facilitate future research on improving the generalization capability of OV-3D models.
@@
A Web Interface for Manual Annotation
We implement a web interface for manual annotation of the visual linguistic aspect (e.g., material), as shown in Figure A. Annotators are shown an interactive 3D mesh of a scene, a list of target objects, and a list of attributes. Users can control the 3D mesh from different viewpoints interactively by rotating, zooming in, zooming out, and panning to observe the scene from various viewpoints. When users select a 3D mesh by clicking the mouse in the scene, the target object will be highlighted and the corresponding object ID and object class will be displayed. The annotation process requires annotators to first select a target object in the 3D mesh (e.g., table of ID 2) and then select a primary attribute that belongs to the target object (e.g., stone). Finally, annotators click the confirm button to submit and store the annotations. Once the selected objects are annotated and confirmed, the corresponding object in the object list will show a check mark symbol. Additionally, to address visual ambiguity issues in 3D object appearance, our annotation process allows users to review the scene’s video sequences, providing contextual visual cues to resolve uncertainties about target objects’ visual attributes during annotation.
B Implementation Details
B.1 Open-Vocabulary 3D Scene Understanding (OV-3D) Baselines
We report implementation details of the OV-3D models (Takmaz et al. 2023; Yin et al. 2024; Yan et al. 2024; Nguyen et al. 2024; Peng et al. 2023; Ding et al. 2023; Yang et al. 2024) as follows:
OpenMask3D. In the class-agnostic mask proposal module, we employ the Mask3D (Schult et al. 2023) architecture trained on the ScanNet200 (Rozenberszki et al. 2022) training set. For 2D mask proposals, we use SAM (Kirillov et al. 2023) with ViT-H as the backbone. We utilize the pre-trained CLIP (Radford et al. 2021) visual encoder of ViT-L/14 at a 336 pixel resolution to extract image features with 768 dimensions. We set the number of queries to 150, following the implementation of OpenMask3D (Takmaz et al. 2023) and Mask3D (Schult et al. 2023), to ensure a sufficient number of mask proposals for the GOV-3D task.
SAI3D. We employ Semantic-SAM (Li et al. 2023) with Swin-L as the backbone to generate 2D mask proposals. The number of queries is set to 150 to ensure sufficient mask proposals for the GOV-3D task.
MaskClustering. We utilize CropFormer (Qi et al. 2023) as a 2D mask predictor. For 2D mask proposals, we use CLIP (Radford et al. 2021) visual encoder of ViT-H/14 to extract image features. We follow MaskClustering (Yan et al. 2024) to adopt the post-processing approach from OVIR-3D (Lu et al. 2023) to refine the output 3D instances. Specifically, we employ the DBSCAN (Ester et al. 1996) algorithm to partition disconnected point clusters.
Open3DIS. We utilize the class-agnostic 3D proposal network ISBNet (Ngo, Hua, and Nguyen 2023) trained on the ScanNet200 (Rozenberszki et al. 2022) training set as 3D proposal. We employ the 2D-Guided-3D Instance Proposal Module in Open3DIS (Nguyen et al. 2024). For 2D mask proposals, we adopt Grounded-SAM (Ren et al. 2024) as 2D segmentor, which incorporates a Swin-T-based Grounding-DINO (Liu et al. 2024) decoder and SAM (Kirillov et al. 2023) with ViT-H as the backbone.
OpenScene. We employ OpenSeg (Ghiasi et al. 2022) for image feature extraction and a 2D-3D ensemble model in OpenScene (Peng et al. 2023). We utilize MinkowskiNet18A (Choy, Gwak, and Savarese 2019) as the 3D backbone during 3D distillation.
PLA. We utilize a model trained on the ScanNet (Dai et al. 2017) partition of B15/N4, where B15/N4 indicates 15 base and 4 novel categories. We adopt a SparseUNet16 architecture based on sparse convolutions UNet (Graham, Engelcke, and Van Der Maaten 2018) as our 3D encoder for semantic segmentation and integrate the CLIP (Radford et al. 2021) text encoder as the final classifier.
RegionPLC. We utilize a model trained on the ScanNet (Dai et al. 2017) partition of B15/N4, where B15/N4 represents 15 base and 4 novel categories. We employ a sparse-convolution-based UNet (Graham, Engelcke, and Van Der Maaten 2018) of SparseUNet16 as the 3D encoder for semantic segmentation, leveraging the CLIP (Radford et al. 2021) text encoder as the final classifier.
@@ B.1 Open-Vocabulary 3D Scene Understanding (OV-3D) Baselines
| OpenMask3D (Takmaz et al. 2023) | 7.7 | 8.2 | 8.7 | 14.2 | 16.3 | 12.0 | 10.0 | 14.7 | 9.7 |
| SAI3D (Yin et al. 2024) | 4.0 | 5.5 | 7.4 | 10.6 | 8.6 | 8.8 | 7.3 | 8.3 | 6.7 |
| MaskClustering (Yan et al. 2024) | 6.0 | 6.2 | 4.1 | 8.0 | 11.7 | 10.3 | 9.3 | 8.1 | 6.8 |
| Open3DIS (Nguyen et al. 2024) | 11.5 | 17.6 | 14.5 | 18.5 | 27.7 | 18.6 | 16.6 | 23.4 | 15.6 |
@@
| OpenMask3D (Takmaz et al. 2023) | 9.6 | 10.7 | 11.2 | 18.1 | 18.4 | 15.2 | 12.8 | 17.8 | 12.2 |
| SAI3D (Yin et al. 2024) | 6.4 | 8.2 | 10.6 | 14.9 | 13.0 | 13.8 | 11.5 | 13.3 | 10.1 |
| MaskClustering (Yan et al. 2024) | 10.3 | 11.5 | 7.2 | 13.5 | 23.1 | 18.7 | 16.1 | 14.0 | 12.0 |
| Open3DIS (Nguyen et al. 2024) | 14.3 | 22.2 | 18.3 | 22.6 | 32.9 | 23.1 | 19.7 | 28.5 | 19.2 |
@@
| OpenMask3D (Takmaz et al. 2023) | 10.8 | 12.4 | 13.9 | 20.4 | 18.9 | 18.0 | 14.4 | 20.5 | 14.1 |
| SAI3D (Yin et al. 2024) | 8.5 | 11.1 | 13.1 | 18.6 | 15.9 | 17.5 | 14.7 | 18.6 | 12.8 |
| MaskClustering (Yan et al. 2024) | 12.3 | 13.5 | 9.4 | 16.9 | 25.1 | 24.1 | 19.0 | 18.1 | 14.6 |
| Open3DIS (Nguyen et al. 2024) | 16.2 | 24.0 | 20.2 | 24.8 | 35.9 | 24.9 | 22.7 | 31.9 | 21.3 |
B.2 Evaluation Protocol
We evaluate the OV-3D baselines on 312 scenes following the validation split of ScanNet200 (Rozenberszki et al. 2022) dataset. Our evaluation includes all 341 attributes across 8 linguistic aspects.
C Additional Experimental Results
C.1 Results of Radar Charts.
Figure B presents radar charts of the main results on our OpenScan benchmark. The charts compare performance across AP, AP50, and AP25 for OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024). Our results demonstrate that Open3DIS achieves the strongest performance across eight linguistic aspects, particularly in terms of AP and AP50. Meanwhile, MaskClustering exhibits competitive performance in AP50 and AP25, with notable strengths in the synonym and requirement aspects.
@@ C Additional Experimental Results
| OpenMask3D (Takmaz et al. 2023) | 19.1 | 12.7 | 28.0 | 26.5 | 9.1 | 0.1 | 41.8 | 16.9 | 23.1 | 10.7 | 18.8 |
| SAI3D (Yin et al. 2024) | 13.1 | 10.3 | 19.2 | 5.6 | 6.5 | 0.1 | 19.7 | 12.8 | 14.9 | 10.4 | 11.3 |
| MaskClustering (Yan et al. 2024) | 12.8 | 18.7 | 26.3 | 10.4 | 6.3 | 0.3 | 25.1 | 13.2 | 4.8 | 3.5 | 12.1 |
| Open3DIS (Nguyen et al. 2024) | 32.9 | 27.1 | 35.4 | 33.0 | 24.6 | 2.4 | 43.1 | 33.8 | 29.9 | 20.9 | 28.3 |
@@
| OpenMask3D (Takmaz et al. 2023) | 23.9 | 16.0 | 30.4 | 30.3 | 11.6 | 0.1 | 44.5 | 20.0 | 29.7 | 15.2 | 22.1 |
| SAI3D (Yin et al. 2024) | 19.6 | 15.7 | 26.9 | 9.5 | 10.2 | 0.1 | 31.8 | 18.4 | 22.6 | 16.2 | 17.1 |
| MaskClustering (Yan et al. 2024) | 23.8 | 31.6 | 41.0 | 19.7 | 12.0 | 0.5 | 37.0 | 23.2 | 9.3 | 7.6 | 20.6 |
| Open3DIS (Nguyen et al. 2024) | 40.9 | 32.7 | 39.0 | 38.6 | 30.7 | 3.5 | 46.3 | 38.8 | 37.6 | 27.8 | 33.6 |
@@
| OpenMask3D (Takmaz et al. 2023) | 27.4 | 19.1 | 32.5 | 33.5 | 13.3 | 0.1 | 47.0 | 21.7 | 34.8 | 21.4 | 25.0 |
| SAI3D (Yin et al. 2024) | 26.4 | 20.5 | 32.2 | 18.5 | 13.7 | 0.2 | 41.0 | 23.4 | 31.4 | 22.0 | 22.9 |
| MaskClustering (Yan et al. 2024) | 31.3 | 38.4 | 49.7 | 23.0 | 16.7 | 0.6 | 41.6 | 28.2 | 16.1 | 10.9 | 25.6 |
| Open3DIS (Nguyen et al. 2024) | 44.7 | 35.3 | 42.6 | 42.2 | 33.5 | 5.1 | 48.3 | 42.4 | 41.8 | 31.6 | 36.7 |